# Sonic4D: Spatial Audio Generation for Immersive 4D Scene Exploration

### v1

@@ 4.1 Experimental Settings
We recruited 25 participants, each of whom evaluated 33 video-audio pairs. All participants completed the experiment in a quiet environment while wearing binaural headphones. To more comprehensively demonstrate our model’s capabilities, we categorized the videos based on custom camera trajectories into static and dynamic viewpoints. Static viewpoints indicate using the original camera view unchanged or only modest transformations of that original view, with the viewpoint remaining fixed throughout; Dynamic viewpoints include translational camera movements, arc-shaped camera trajectories, and push-in/pull-back motions.
For data processing, we first excluded participants whose ratings exhibited low Pearson correlation with the overall preference distribution. We then computed and reported the preference percentages and mean opinion scores (MOS) across all viewpoints, as well as separately for static and dynamic viewpoints.
4.2 Quantitative Comparisons
Metric All Static viewpoints Dynamic viewpoints
@@ 4.2 Quantitative Comparisons
MOS-AVSC
   MMAudio cheng2024taming 2.418 2.493 2.363
@@
Table 1: Subjective evaluation results for Sonic4D vs. MMAudio cheng2024taming .
We present the results of user study evaluation comparing the spatial audio rendered by our method (Sonic4D) against the non-spatialized mono audio from MMAudio cheng2024taming . As shown in Table 1, Sonic4D achieved a strong overall preference rate of 89.03%, indicating that the spatial audio generated by Sonic4D exhibits a significant perceptual difference in spatiality compared to the original mono audio, thus demonstrating the effectiveness of our model. This trend held consistently across both static viewpoints (87.59%) and dynamic viewpoints (90.10%), demonstrating the robustness of our approach under varying camera trajectories.
In terms of Mean Opinion Scores (MOS), our method significantly outperforms the baseline in both evaluation dimensions. For Spatial Localization Accuracy (MOS-SLA), Sonic4D achieved a mean score of 4.013, compared to MMAudio cheng2024taming ’s 2.322. This suggests that participants were clearly able to perceive accurate and stable sound source positions when listening to our spatialized audio. A similar advantage is observed for Audio–Visual Spatial Congruence (MOS-AVSC), where Sonic4D scored 3.977 versus 2.418 for MMAudio cheng2024taming , indicating a much stronger alignment between auditory motion cues and on-screen visual dynamics.
Overall, these results validate that Sonic4D produces compelling and perceptually convincing spatial audio that generalizes well across diverse viewing conditions. Participants not only preferred the spatialized version, but also rated it higher in terms of spatial realism and audio-visual coherence, supporting the effectiveness of our approach in rendering viewpoint-consistent spatial audio.
4.3 Qualitative Comparisons
@@
In case 1, the helicopter first moves leftward and then rightward across the frame, resulting in a perceived spatial effect where the sound shifts from being centered, to the left, and then to the right. By comparing Regions 1 and 2 with Regions 3 and 4 in the spatial audio waveform, we observe that the left channel initially has greater amplitude than the right. As the helicopter moves rightward, the left channel amplitude rapidly decreases and eventually falls below the right, aligning well with the expected spatial perception of the moving object. In case 2, the camera starts on the right side of a ukulele player and orbits to the left. The left-right disparity in Regions 1 and 2 , as well as the temporal progression within those regions, closely mirrors the trajectory, indicating strong spatial tracking in the generated audio. Case 3 demonstrates that the same source video can yield notably different spatial audio depending on the camera motion, validating the view-dependence of our model. In video 1, as the camera pans rightward, the fountain becomes closer to the left ear, leading to increased amplitude in the left channel (Regions 1,2 ). In video 2, the camera gradually pulls out, producing a fading splash sound that matches the increasing distance (Regions 3,4,5,6 ).
@@ 5 Conclusions
6 Limitations
Although our method can generate spatial audio that follows the camera viewpoint with clear directional cues, it still has several potential limitations:
  • While we model the dynamic trajectories of both the sound source and the listener using physics-based methods, the quality of the synthesized binaural audio may still be limited due to the lack of fine-grained acoustic simulation, such as incorporating Head-Related Transfer Functions (HRTF) or complex wave-based propagation models.
  • For simplicity, we constrain our setup to a single sound source per video. However, thanks to the flexibility of our framework, multi-source localization and generation can be supported by modifying the prompts provided to GroundingGPT and MMAudio accordingly.
  • Our use of monocular depth–derived point clouds lacks detailed geometric modeling, which may limit the physical accuracy of spatial audio propagation in complex environments.
Appendix A Details of User Study
A.1 Experiment Setup and Procedure
The user study consists of three main stages:
Step 1: Dynamic Scene and Monaural Audio Generation. In the first stage of our pipeline, as described in the Dynamic Scene and Monaural Audio Generation section, we used TrajectoryCrafter [54] to predict dynamic point clouds for 4D scene construction. Since TrajectoryCrafter [54] uses the relative transformation between the user-specified camera trajectory and the source camera as the basis for novel view rendering and requires a stabilized source video, we selected only source videos that meet this requirement. To generate these videos, we manually composed prompts for commercial text-to-video models (e.g., Kling [60], Dreamina [61], Vidu [62]), and for each sample, we designed a custom camera trajectory to render the target view. We then leveraged Sonic4D to generate both the novel view video and the corresponding spatialized audio.
Step 2: Interface Setup and Counterbalancing. As illustrated in Figure4, we created a local web-based interface to facilitate user evaluation. The landing page provided clear instructions on the task and rating criteria as shown in 4(a). We created 33 question pages for participants to complete. To ensure counterbalancing, we randomized the order of all audio-video pairs as shown in 4(b), as well as the internal ordering of each pair (left/right assignment), to mitigate any potential order effects or positional biases. Participants were asked to rate each video on the following two aspects using a 5-point Likert scale. The detailed rating criteria are as follows:
  • Spatial Localization Accuracy:
    • 5 – The position of the subject or camera can be clearly perceived from the audio.
    • 4 – The direction or positional changes can generally be perceived.
    • 3 – Some spatial positioning is conveyed through the audio.
    • 2 – Audio provides vague or weak spatial cues.
    • 1 – Little to no sense of spatial position from the audio.
  • Audio–Visual Spatial Congruence:
    • 5 – The direction indicated by audio cues (e.g., volume changes, binaural differences) perfectly matches the visual motion.
    • 4 – The audio directionality mostly matches the visuals.
    • 3 – The audio and visual alignment is partially consistent.
    • 2 – The audio and visuals are misaligned or direction is hard to perceive.
    • 1 – The spatial audio cues conflict with or fail to reflect the visuals.
(a) User study instruction page outlining participation guidelines and evaluation criteria.
(b) Example of a user study question page, displaying paired videos and corresponding rating options.
Figure 4: Screenshots of the user study interface.
Step 3: Data Collection and Filtering. After collecting all participant responses regarding spatial audio preference and ratings, we computed the Pearson correlation between each participant’s responses and the overall average responses. Specifically, for each participant $i$, we encode their answers on all preference questions as a numerical vector $\mathbf{a}^{(i)}$, where each answer is mapped to a scalar:
$a^{(i)}_{j}=\begin{cases}1,&\text{if the participant preferred Video 1}\\ 2,&\text{if the participant preferred Video 2}\\ 0,&\text{if the participant selected ``Almost same''}\end{cases}$ (12)
Here, $a^{(i)}_{j}$ denotes the response of the $i$-th participant to the $j$-th question, where $j=1,2,\dots,M$. The full response vector for participant $i$ is:
$\mathbf{a}^{(i)}=[a^{(i)}_{1},a^{(i)}_{2},\dots,a^{(i)}_{M}]$
We then compute the mean preference vector across all $N$ participants:
$\bar{\mathbf{a}}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{a}^{(i)}$ (13)
Next, the Pearson correlation coefficient between each participant’s response vector and the mean vector is computed as:
$r^{(i)}=\frac{\sum_{j=1}^{M}\left(a^{(i)}_{j}-\bar{a}^{(i)}\right)\left(\bar{a}_{j}-\bar{\bar{a}}\right)}{\sqrt{\sum_{j=1}^{M}\left(a^{(i)}_{j}-\bar{a}^{(i)}\right)^{2}}\sqrt{\sum_{j=1}^{M}\left(\bar{a}_{j}-\bar{\bar{a}}\right)^{2}}}$ (14)
where $M$ is the number of preference questions, $\bar{a}^{(i)}$ is the mean of participant $i$’s answers, and $\bar{\bar{a}}$ is the mean of the overall mean vector $\bar{\mathbf{a}}$.
We exclude participant $i$ if $r^{(i)}<0.4$, considering such responses to be poorly correlated with the overall trend. In the end, we retained 22 valid participant responses, and calculated the final Mean Opinion Scores (MOS) based on these.
A.2 Experimental Video Categories
To analyze the impact of different viewpoint conditions on spatial audio perception, we categorized the evaluation videos into two major groups: static and dynamic viewpoints.
A video is considered to have a static viewpoint if the camera position is fixed throughout the clip, or undergoes only minor transformations of original camera pose but still remains fixed. In contrast, a dynamic viewpoint involves explicit camera motion, such as linear translation, arc-shaped trajectories around a subject, or push-in/pull-back zooming effects that create continuous changes in perspective.
To ensure a balanced evaluation, we selected a diverse set of examples from each category, as summarized in Table2. The dataset includes a roughly balanced distribution of static and dynamic scenarios.
View Type Subtype Number of Samples
Static Raw viewpoint (static sound source) 3
Raw viewpoint (dynamic sound source) 7
Fixed camera (transformation) 4
Total (Static) 14
Dynamic Translational movement 5
Arc-shaped trajectory 7
Push-in / Pull-back 7
Total (Dynamic) 19
Table 2: Distribution of evaluation videos by viewpoint category
Appendix B Details of Physics-Driven Spatial Audio Synthesis
Room Impulse Response (RIR) plays a crucial role in spatial audio rendering by modeling how sound propagates and reflects within an environment. Accurate RIR simulation enables the synthesis of physically plausible spatial cues, enhancing the realism of the generated audio experience.
In our pipeline, we adopt gpuRIR [59] as the simulator to generate synthetic RIRs. To reflect varying acoustic environments, we predefine four types of room size categories with randomly sampled dimensions in meters, as shown below:
  • Small: Length and width $\in$ [3, 8], height $\in$ [2.5, 4]
  • Medium: Length and width $\in$ [8, 15], height $\in$ [3, 6]
  • Large: Length and width $\in$ [15, 30], height $\in$ [5, 10]
  • Outdoor: Length and width $\in$ [100, 200], height $\in$ [50, 100]
For each category, the length, width, and height are randomly sampled within the corresponding range. Given the source and receiver trajectories derived in Stage 2, we compute the minimal bounding cuboid that encapsulates both trajectories. This cuboid is then placed at the center of the room. If the bounding box exceeds the selected room size, it is uniformly scaled down to fit the lower bounds of the selected category.
To ensure smooth transitions and minimize auditory artifacts, we perform temporal upsampling on the trajectories before simulation. Specifically, we use a high-resolution sampling rate of $n=44100$ points per trajectory.
For microphone orientation, we support two polar patterns: omnidirectional (omni) and cardioid (card). In the cardioid setting, the microphone orientation at each point is aligned to face outward from the center plane of the rendering space.
Finally, the reverberation time ($T_{60}$) is randomly sampled from a uniform distribution in the range (0.3, 0.6) seconds to simulate varying acoustic characteristics across environments.

### v3

@@ Methods
To more comprehensively demonstrate our model’s capabilities, we categorized the videos based on custom camera trajectories into static and dynamic viewpoints. Static viewpoints indicate using the original camera view unchanged or only modest transformations of that original view, with the viewpoint remaining fixed throughout; Dynamic viewpoints include translational camera movements, arc‑shaped camera trajectories, and push‑in/pull‑back motions.
Quantitative Comparisons
SELD Evaluation.
Table 1 summarizes the SELD evaluation results on our curated STARSS23 subset. Our method attains a relatively higher F-score (4.0%) than ViSAGe, while reducing DOA angular error to 18.6° and distance errors to 75.20 (absolute) and 0.44 (relative). These results demonstrate that Sonic4D produces more accurate spatial audio cues than purely learning-based methods, highlighting its ability to capture spatial information in a zero-shot setting.
[…]
Ground truth 6.5% 7.2 23.67 0.15
ViSAGe 0.4% 31.5 101.18 0.49
Sonic4D(ours) 4.0% 18.6 75.20 0.44
Table 1: SELD evaluation results on the curated STARSS23 subset for Sonic4D vs. ViSAGe (Kim et al. 2025).
User Study.
We present the results of user study evaluation comparing the spatial audio rendered by our method (Sonic4D) against the non-spatialized mono audio from MMAudio (Cheng et al. 2024). As shown in Table 2, Sonic4D achieved a strong overall preference rate of 89.03%, indicating that the spatial audio generated by Sonic4D exhibits a significant perceptual difference in spatiality compared to the original mono audio, thus demonstrating the effectiveness of our model. In terms of Mean Opinion Scores (MOS), our method outperforms the baseline in both evaluation dimensions. For Spatial Localization Accuracy, Sonic4D achieved a mean score of 4.013, compared to MMAudio (Cheng et al. 2024)’s 2.322. This suggests that participants were clearly able to perceive accurate and stable sound source positions when listening to our spatialized audio. A similar advantage is observed for Audio–Visual Spatial Alignment, where Sonic4D scored 3.977 versus 2.418 for MMAudio (Cheng et al. 2024), indicating a much stronger alignment between auditory motion cues and on-screen visual dynamics.
Metric All Static Dynamic
@@
MOS-AVSA
   MMAudio 2.418 2.493 2.363
@@
Table 2: Subjective evaluation results for Sonic4D vs. MMAudio (Cheng et al. 2024).
Qualitative Comparisons
@@
In Case 1, as the helicopter moves left then right, the spatial audio waveform reflects this motion: early segments (Regions 1 and 2) show the left channel dominant, whereas later segments (Regions 3 and 4) show the right channel surpassing the left corresponding to the helicopter’s trajectory, aligning well with the expected spatial perception of the moving object. In case 2, the camera starts on the right side of a ukulele player and orbits to the left. The left-right disparity in Regions 1 and 2 (Yellow), as well as the temporal progression within those regions, closely mirrors the trajectory, indicating strong spatial tracking in the generated audio. Case 3 demonstrates that the same source video can yield notably different spatial audio depending on the camera motion, validating the view-dependence of our model. In video 1, as the camera pans rightward, the fountain becomes closer to the left ear, leading to increased amplitude in the left channel (Regions 1, 2, Green). In video 2, the camera gradually pulls out, producing a fading splash sound that matches the increasing distance (Regions 3, 4, 5, 6, Green).